OpenStudy (anonymous):

Sequences and series Question...Please help!!! A sequence is defined by Un = 3n - 2. a. Prove that the sequence is arithmetic. Hint: Find Un+1 - Un b. Find U1 and d. c. Find the 57th term. d. What is the least term of the sequence which is greater than 450?

hartnn (hartnn):

which part you have doubt with ?

OpenStudy (anonymous):

Part i dont understand, all the others i know how to do.

OpenStudy (anonymous):

a. sorry

hartnn (hartnn):

A sequence is arithmetic if Un+1 - Un is constant.

hartnn (hartnn):

what you got \(U_{n+1}\)as ?? could you find it ?

OpenStudy (anonymous):

3n+1

hartnn (hartnn):

so,U n+1 - Un =....??

OpenStudy (anonymous):

that equals 3

OpenStudy (anonymous):

what does the three mean?

hartnn (hartnn):

since its constant, the sequence is arithmetic. that Un+1 - Un is actually d = the common difference

hartnn (hartnn):

A sequence is arithmetic if d = the common difference=Un+1 - Un is constant.

OpenStudy (anonymous):

oh... ok i get it now. thnx a lot. i did figure out that d is the same as Un+1 -Un

hartnn (hartnn):

welcome ^_^